OpenStudy (anonymous):

A roll of wallpaper covers an area of 28ft squared. If the roll is 1 17/24 ft wide, how long is the roll? I have been working on this problem for over an how, I need to see it step by step.

OpenStudy (anonymous):

Well a roll of wallpaper is just a large rectangular piece of wallpaper. So you know the width and the area, so you can find the length. L * W = A so.... L= A/W

OpenStudy (anonymous):

the width is 1 17/24 ft. the are is 28x28=784, correct

OpenStudy (anonymous):

You have the area and the width, you can find the length by using the fact that Area=lw and solve for l to give you l=Area/width to get your answer...

OpenStudy (dumbcow):

trouble, you don't need to square the 28....they have already given you the area which is 28 also get rid of the mixed number and make it a fraction -->41/24 divide: 28/(41/24) = 28*(24/41) = 672/41 =16.39

OpenStudy (anonymous):

I did exactly what you did. My answer is 16 20/41.

OpenStudy (dumbcow):

16 20/41 = 676/41